Fermi National Laboratory


Accelerator Update

Friday November 14
The day shift began with Operations preparing for a power outage to restore power to the Master Substation.

The power outage began at 10 AM and lasted ten minutes. Operators and experts began restoring accelerator systems.

At 12:40 PM, Linac experts reported that Klystron 4 needed replacing before they could startup. They began work immediately.

Operations received permission to send beam to the Recycler at 1:24 PM and permission run the Tevatron (TeV) at 2:20 PM.

Both the TeV and Main Injector (MI) RF was up and running by 3 PM.

At 4:43 PM, a Linac expert reported that a Klystron water pump failed due to an electrical wiring problem. A system expert was notified. See picture below.

Technicians completed their repairs on KRF4 at 7:30 PM and experts established beam through Linac just a few minutes later. They had the beamline tuned by 8:15 PM and beam to the Booster dump by 8:30 PM.

Cryo system experts reported that CHL was cold at 8:43 PM. They began filling the transfer lines to cool down the TeV.

Saturday November 15
The midnight shift began with experts tuning Booster and MI.

By 1 AM, MI experts were ready to send beam to the Antiproton Source (Pbar).

Experts established protons to the Pbar Accumulator at 3:14 AM and had reverse protons to the Debuncher by 5:20 AM.

At 10:55 AM, a Linac RF station (KRF1) tripped off. Operators tried to reset the station, but it wouldn't stay up. A Linac tech came in and made adjustments. He had the station operating normally by 1:10 PM.

The Collider Detector Facility powered their solenoid at 2:40 PM.

Around 7 PM, Operations began sending beam to MiniBooNE.

Sunday November 16
The midnight shift began with experts tuning and with MiniBooNE taking beam.

A 12:50 AM, MI had the first of eight or nine trips due to vacuum bursts over a forty-five minute period.

Pbar began stacking at 1:40 AM.

At 3:15 PM, cryo techs reported a vacuum problem in the TeV sector E2 area. Vacuum techs came in and accessed the TeV at 6 PM. They attached another vacuum pump, which resolved the problem temporarily.

The TeV went cold at 6 PM. TeV turn on began and successfully ramped to 980 GeV at 9:30 PM.

Monday November 17
The midnight shift began with beam studies in most areas, with TeV turn on progressing, and with MiniBooNE taking beam.

TeV experts established beam to the TeV at 12:45 AM and established circulating beam by 2:30 AM.

Two Booster RF stations (BRF5 & 6) tripped due to a vacuum burst at 5:30 AM. They both reset okay.

Recyler experts took beam from MI and successfully circulated it.

Today's Plans
Continue tuning accelerators and checking out systems.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 11/17/2003   email Fermilab